# SceneVerse: Scaling 3D Vision-Language Learning for Grounded Scene Understanding

### v1

@@ 4 Grounded Pre-training for Scenes
Figure 3: Overview of our proposed GPS model. We leverage contrastive alignment in three levels $\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{scene}}$, and $\mathcal{L}_{\text{ref}}$ as well as a masked language modeling objective $\mathcal{L}_{\text{MLM}}$ for model learning.
In this section, we introduce GPS, an efficient transformer-based model trained with multi-level contrastive losses for aligning 3D scenes and texts. As shown in Fig. 3, we echo the language descriptions collected at different levels to form scene-language pairs at both object-level, referral-object-level, and scene-level for contrastive objectives in GPS. We describe the design of each level in the following sections.
@@ 4.1 Object-level Grounding
$\mathcal{L}_{\text{obj}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(p,r)\right)}}+\right.$ (1)
$\left.\ \ \ \log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(r,q)\right)}}\right),$
@@ 4.2 Scene-level Grounding
$\mathcal{L}_{\text{scene}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(p,r)\right)}}+\right.$ (2)
$\left.\ \ \ \log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(r,q)\right)}}\right),$
@@ 4.3 Referral-object-level Grounding
To learn the multi-level alignment between 3D scenes and language, we first train the point cloud encoder with object-level grounding objects to obtain a good feature initialization for grounding objects in scenes. During the scene grounding stage, we train our inter- and intra-scene objectives together with a mask language modeling loss $\mathcal{L}_{\text{MLM}}$ over the inputted object-referral texts to tune the parameters within the language encoder and self-attention transformer. Above all, the learning of GPS could be summarized as optimizing the following objective:
@@ 5 Experiments
In this section, we present the evaluation results addressing the following questions:
  • How effective is the data scaling in SceneVerse for 3D visual grounding? Does the scale-up work for general pre-training based 3D-VL models?
  • How well is the GPS pre-training pipeline? Does it exhibit similar properties of 2D-VL models in 3D-VL tasks?
  • What potentials are offered by SceneVerse and GPS for future research? What is missing?
In the following sections, we describe in detail about model performance regarding these key topics. Due to the page limit, we direct readers to the Appendices B and C for implementation details and more experimental analyses.
@@
Table 2: 3D Visual Grounding results on Nr3D, Sr3D, and ScanRefer. We use “direct” for our model trained on SceneVerse with no additional fine-tune head, and “fine-tune” for the data-specific fine-tuned version of our model. We highlight the best results in bold.
@@
Settings
We evaluate our model on three commonly-used datasets for 3D visual grounding: ScanRefer [16], Nr3D, and Sr3D [1]. For Nr3D and Sr3D, we follow Achlioptas et al. [1] and report the grounding accuracies of models using ground-truth object masks. For ScanRefer, we follow Zhu et al. [98] and use Mask3D [72] to generate object proposals. Results are reported as $\text{Acc}@0.5$ to evaluate the correctness of predictions whose object bounding boxes overlap the ground truth with IoU $>0.5$. For comparisons, we compare with existing baselines by providing the results of pre-trained GPS and dataset-specific fine-tuned GPS. Please see more details in the Appendix C.
Results and Analyses
@@ 5.2 Zero-Shot Transfer
To better evaluate the effectiveness of both the SceneVerse data and the GPS model, we further perform zero-shot transfer experiments to test the models’ capability in 4 benchmarks, ScanRefer, Sr3D, Nr3D, and SceneVerse-val. We create SceneVerse-val using $8.5K$ annotated object referrals of $271$ scenes in MultiScan, and randomly split the scenes following a 4:1 train / test split for creating the held-out test set. We mainly consider 2 specific transfer settings in our experiments: (i) zero-shot: models trained by removing all the scenes from the target dataset, tested on held-out unseen scenes, and (ii) zero-shot text: Models trained on data that include the 3D scenes from training set of the target dataset, yet tested exclusively with unseen scene-text distribution. Specifically, for the zero-shot text setting, we use the generated texts in SceneVerse as fine-tuning sources for the zero-shot model. We mainly compare our model against a recent pre-training-based model 3D-VisTA. See more details on experimental setting and implementation in the Appendix C.
Table 3: Zero-shot transfer results on established benchmarks.
Method Nr3D Sr3D ScanRefer@0.25 ScanRefer@0.5
@@
Table 4: Zero-shot transfer on SceneVerse-val. We evaluate models following settings in Nr3D/Sr3D using GT object proposals.
@@
We present the results of zero-shot transfer experiments in Tab. 3 and Tab. 4 with the following key observations:
@@
  • SceneVerse dataset substantially enhances 3D-VL grounding capabilities through zero-shot transfer, especially when provided with relatively limited training data, i.e., SceneVerse-val. As demonstrated in  Tab. 4, there is a significantly improved performance when comparing models trained on SceneVerse in a zero-shot manner to those trained from scratch. This indicates that SceneVerse can effectively capture knowledge for general 3D scene grounding. Consequently, this underscores its potential as a go-to pre-training dataset for 3D-VL tasks.
  • The impact of our extensive collection and scalable generation of scene-text pairs is further evidenced by the results in the zero-shot text setting. Notably, as shown in Tab. 3, these automatically generated scene-text pairs supply ample knowledge for comprehending the scene distribution. This contributes significantly to the substantial improvement over the zero-shot performance.
Figure 4: Model performance v.s. data scale. Models consistently improve in both the pre-train and zero-shot transfer settings on ScanRefer and SceneVerse-val with data scaling-up.
5.3 Ablative Studies and Discussions
In this section, we present ablative studies focused primarily on the data collected in SceneVerse. Our aim is to elucidate the effects of data scaling more clearly. For discussions regarding ablation studies on model architecture, readers are referred to the Appendix D. The following points are specifically discussed in this section.
@@ 5.3 Ablative Studies and Discussions
We conduct ablation studies over the amount of data used while pre-training GPS. We consider the model trained with $\frac{1}{8}$, $\frac{1}{4}$, $\frac{1}{2}$ of SceneVerse to show the effectiveness of data-scaling on model performance in the pre-train and zero-shot transfer settings in ScanRefer and SceneVerse-val. As shown in Fig. 4, we observe consistent performance improvement over the increase of data scale for both settings. We provide additional experiments in the Appendix D to show that such scaling effect is not only beneficial for 3D-VL grounding but also for other 3D tasks like semantic segmentation [72, 85].
How is the generated data compared with human-annotated data?
We assess the performance of models trained using various scene-text sources, specifically focusing on their performance in the ScanRefer dataset without additional fine-tuning. As shown in Tab. 5, models trained with our template-based generated texts and Large Language Model (LLM)-refined texts show significant improvements over models trained solely on ScanRefer. More importantly, these variants of our model already achieve state-of-the-art results compared with previous baselines. This indicates the effectiveness of our text-generation pipeline. Finally, we observe that adding human-annotated data is still beneficial for model performance. However, the improvement is relatively marginal over models trained on our generated data.
What is the role of the synthetic scenes in this scale-up process?
With synthetic data providing large-scale and diverse scene data for 3D-VL tasks, we evaluate the models’ domain transfer (Sim2Real) capability. Specifically, we compare models trained on all real scenes in SceneVerse against models trained exclusively on two synthetic subsets of SceneVerse, i.e., Structured3D and ProcTHOR. As shown in Tab. 6, models trained on synthetic subsets demonstrate remarkable performance on their corresponding test sets while suffering when transferred to real or other synthetic scenes. In contrast, the model trained on real scene-text pairs exhibits less severe performance drops when generalizing to synthetic scenes. This result affirms the domain gap between real and synthetic scenes in 3D-VL grounding and shows that a simple scale-up in the amount of scenes is insufficient when the scene naturalness can not be guaranteed. Considering the scalability of our quality-ensured language generation and also the scaling effect shown in our experiments, the rate-determining step for further scaling-up 3D-VL comes to the collection of diverse, high-quality, and realistic scenes that capture natural 3D scene distributions.
Table 5: Ablation on scene-text pair types used in training. We report model results on ScanRefer with no additional finetuning.
@@
Table 6: Cross domain transfer results of models learned in real and synthetic datasets without additional fine-tuning. “S3D” stands for Structured3D.
@@ 6 Conclusion
In this work, we scale up 3D-VL in the context of grounded scene understanding. We introduce SceneVerse, a million-scale 3D-VL dataset encompassing various scenes and multi-level scene descriptions sourced from both human annotation and our proposed scene-text generation approach. Utilizing SceneVerse, we propose Grounded Pre-training for Scenes, a model trained with multi-level scene-language contrastive alignment over the data collected. Through extensive experiments, we show that GPS achieves state-of-the-art results on all existing 3D-VL grounding tasks. We further conduct zero-shot transfer experiments to show the improved generalization performances of GPS trained on SceneVerse compared with previous baselines. We hope our efforts and successful scale-up attempts in SceneVerse could pave the way for a new research paradigm in 3D-VL.
7 Acknowledgement
The authors thank Yaowei Zhang from BIGAI for designing the result visualization framework, Jiangyong Huang and Xiongkun Linghu from BIGAI for suggestions on data generation and refinement, and colleagues from BIGAI for their helpful discussions and suggestions.
@@ 7 Acknowledgement
In Appendix A, we introduce more details of SceneVerse, including the 3D scene preprocessing, scene graph construction and automatic language generation. Appendix B presents more model and implementation details. Appendix C include a more in-depth summary of the settings and implementations for the experiments in the main paper, as well as the ablative study and additional experiments on semantic segmentation to demonstrate the benefits of SceneVerse.
Appendix A SceneVerse
A.1 3D Scenes
@@ Appendix A SceneVerse
In order to mitigate the data disparities arising from diverse capture devices across various data sources, we subsample each point cloud to a maximum of $240,000$ points. Each point cloud then undergoes a transformation centered on the central point on the floor, followed by rotation to align the room layout with the axis following the approach by Chen et al. [18].
@@
A.2 3D Scene Graph Construction
@@ A.2.1 Relationships
|||| supported by
In-contact vertical embedded into
placed in
|||| inside
hanging on
affixed on
mounted on
Non-contact vertical above
higher than
below
@@
near(far) to the left of
near(far) to the right of
is behind
Horizontal is in front of
close to
adjacent to
besides
next to
Multi-object between
aligned
@@
A.2.2 Scene Graph Construction
Due to the inherent noise and incompleteness in the point cloud representation, automatically extracting precise and comprehensive relationships from the point clouds is a non-trivial task. Below we detail our 3D scene graph construction process, as outlined in Alg. 1.
@@ A.2.2 Scene Graph Construction
In our constructed 3D scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the nodes $\mathcal{V}$ comprises $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level. The hierarchies are determined by the support relationship; for instance, objects supported by the floor constitute $\mathcal{V}_{0}$, while objects supported by the table will form $\mathcal{V}_{1}$, etc. Note that edges originating from one node $v\in\mathcal{V}_{k}$ may only terminate in nearby hierarchies $\mathcal{V}_{k}\cup\mathcal{V}_{k+1}\cup\mathcal{V}_{k+1}$. In other words, edges in the scene graph exclusively connect nodes within the same hierarchical level, or one level higher or lower.
@@
2:      Create node $v_{i}\in\mathcal{V}$ using the centroid $p_{i}$ and bounding box size $b_{i}$ of object point cloud $P_{i}$
@@
7:         Add in-contact vertical relationship triplets $(v_{i},v_{j},e_{i,j})$ with $\mathrm{RelsType}_{v}$ to $\mathcal{G}$
@@
12:            Add non-contact vertical relationship triplets $(v_{i},v_{k},e_{i,k})$ with $\mathrm{RelsType}_{v}$ to $\mathcal{G}$
@@
17:      let ${\{v_{i_{1}},v_{i_{2}},...,v_{i_{N}}\}}$ be the $N$ different nodes with the same in-contact vertical parent node $v_{i}$

### v3

@@ 4 Grounded Pre-training for Scenes
Figure 3: Overview of GPS model. We use contrastive alignment at three levels $\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{scene}}$, and $\mathcal{L}_{\text{ref}}$ and a masked language modeling objective $\mathcal{L}_{\text{MLM}}$ for model learning.
In this section, we introduce GPS, an efficient transformer-based model trained with multi-level contrastive losses for aligning 3D scenes and texts. As shown in Fig. 3, we echo the language descriptions collected at different granularities to form contrastive objectives at both object-level, referral-object-level, and scene-level in GPS. We describe the design of each level in the following sections.
@@ 4.1 Object-level Grounding
$\mathcal{L}_{\text{obj}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(p,r)\right)}}+\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(r,q)\right)}}\right),$ (1)
@@ 4.2 Scene-level Grounding
$\mathcal{L}_{\text{scene}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(p,r)\right)}}+\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(r,q)\right)}}\right),$ (2)
@@ 4.3 Referral-object-level Grounding
To learn the multi-level alignment between 3D scenes and language, we first train the point cloud encoder with an object-level grounding objective to obtain a good feature initialization for grounding objects in scenes. During the scene grounding stage, we train our inter- and intra-scene objectives together with a masked language modeling loss $\mathcal{L}_{\text{MLM}}$ over the inputted object-referral texts to tune the parameters within the language encoder and self-attention transformer. Above all, the learning of GPS could be summarized as optimizing:
@@ 5 Experiments
In this section, we present experimental results addressing the following questions:
  1. 1.
    How effective is the data scaling in SceneVerse for 3D visual grounding? Does the scale-up benefit common 3D-VL tasks (e.g., 3D question answering, open-vocabulary 3D semantic segmentation) and pre-training-based models?
  2. 2.
    How well is the GPS pre-training pipeline for 3D-VL tasks? Does it exhibit similar properties of 2D-VL models in 3D-VL tasks?
  3. 3.
    What is offered by SceneVerse and GPS and what is missing?
In the following sections, we describe in detail the model performance regarding these key topics. Due to the page limit, we direct readers to the supplementary for implementation details, qualitative results, and more experimental analyses.
@@
Settings
We evaluate our model on three commonly-used datasets for 3D visual grounding: ScanRefer [16], Nr3D, and Sr3D [1]. For Nr3D and Sr3D, we follow Achlioptas et al. [1] and report the grounding accuracies of models using ground-truth object masks. For ScanRefer, we follow Zhu et al. [112] and use Mask3D [83] to generate object proposals. Results are reported as $\text{Acc}@0.5$ to evaluate the correctness of predictions whose object bounding boxes overlap the ground truth with IoU $>0.5$. For comparisons, we compare with existing baselines by providing the results of pre-trained GPS and dataset-specific fine-tuned GPS. Please see more details in the supplementary.
Results and Analyses
Table 2: 3D visual grounding results on Nr3D, Sr3D, and ScanRefer. We use “pre-train” for our model trained on SceneVerse w/o additional fine-tuning, and “fine-tune” for its data-specific fine-tuned version. Best results are highlighted in bold.
@@ 5.2 Zero-Shot Transfer
To better evaluate the effectiveness of both the SceneVerse data and the GPS model, we further perform zero-shot transfer experiments to test the models’ capability in 4 benchmarks, ScanRefer, Sr3D, Nr3D, and SceneVerse-val. We create SceneVerse-val using $8.5K$ annotated object referrals of $271$ scenes in MultiScan, and randomly split the scenes following a 4:1 train / test split for creating the held-out test set. We mainly consider 2 specific transfer settings in our experiments: (i) zero-shot: models trained by removing all the scenes from the target dataset, tested on held-out unseen scenes, and (ii) zero-shot text: Models trained on data that include the training set of scenes from the target dataset, yet tested exclusively with unseen scene-text distribution. Specifically, for the zero-shot text setting, we use the generated texts in SceneVerse as fine-tuning sources for the zero-shot model. We mainly compare our model against a recent pre-training-based model 3D-VisTA. See more details on experimental setting and implementation in the supplementary.
Table 3: Zero-shot transfer on existing benchmarks. “SR” stands for ScanRefer.
Method Nr3D Sr3D SR@0.25 SR@0.5
@@
Table 4: Zero-shot transfer on SceneVerse-val. Evaluation uses GT object proposals following Nr3D/Sr3D.
@@
We present the results of zero-shot transfer experiments in Sec. 5.2 and Sec. 5.2 with the following key observations:
[…]
@@
[…]
    SceneVerse dataset substantially enhances 3D-VL grounding capabilities through zero-shot transfer, especially when provided with relatively limited training data, i.e., SceneVerse-val. As demonstrated in  Sec. 5.2, there is a significantly improved performance when comparing models trained on SceneVerse in a zero-shot manner to those trained from scratch. This indicates that SceneVerse can effectively capture knowledge for general 3D scene grounding. Consequently, this underscores its potential as a go-to pre-training dataset for 3D-VL tasks.
[…]
    The impact of our extensive collection and scalable generation of scene-text pairs is further evidenced by the results in the zero-shot text setting. Notably, as shown in Sec. 5.2, these automatically generated scene-text pairs supply ample knowledge for comprehending the scene distribution. This contributes significantly to the substantial improvement over the zero-shot performance.
Table 5: 3D question answering results on ScanQA and SQA3D. We report EM@1 score on ScanQA and SQA3D evaluation sets.
Model ScanQA SQA3D
val w/obj w/o obj
ScanRefer+MCAN [5] 18.6 20.6 19.0 -
ScanQA [5] 20.3 23.5 20.9 46.6
SQA3D [67] - - - 47.2
3D-VisTA [112] 22.4 27.0 23.0 48.5
3D-LLM [42] 20.5 19.1 - -
Ours 22.7 25.0 23.5 49.9
Table 6: Exisiting 3D backbones pre-trained on SceneVerse for open-vocabulary 3D semantic segmentation on ScanNet. “SPUNet” denotes SparseUNet proposed in [97].
Model Network mIoU $\Delta$ mAcc $\Delta$
OpenScene [76] SPUNet16 57.2 - 69.9 -
PLA [31] SPUNet16 17.7 - 33.5 -
RegionPLC [97] SPUNet16 56.9 - 75.6 -
RegionPLC+SceneVerse SPUNet16 58.2 +1.7% 77.3 +2.2%
OpenScene [76] SPUNet32 57.8 - 70.3 -
PLA [31] SPUNet32 19.1 - 41.5 -
RegionPLC [97] SPUNet32 59.6 - 77.5 -
RegionPLC+SceneVerse SPUNet32 61.0 +2.3% 79.7 +2.8%
5.3 Additional 3D-VL Tasks
Settings
We evaluate the effectiveness of GPS and SceneVerse on additional 3D-VL tasks: (i) 3D question answering (3D-QA) on ScanQA [5] and SQA3D [67], and (ii) open-vocabulary 3D semantic segmentation (OV-Seg) on ScanNet.
[…]
    In the 3D-QA task, we follow Zhu et al. [112] and evaluate models over the exact match metic (EM@1) on the validation and test sets of ScanQA, as well as the test set of SQA3D. We pre-train GPS on SceneVerse and fine-tune the model on the 3D-QA dataset to compare with state-of-the-art models.
[…]
    In the OV-Seg task, as GPS builds upon an object-centric design and thus is not directly applicable to semantic segmentation, we consider testing the effectiveness of SceneVerse on improving existing 3D models. Specifically, we follow the open-vocabulary semantic segmentation settings proposed by Yang et al. [97] and report the mIoU and mAcc score. We compare with existing works by pre-training the RegionPLC [97] model on SceneVerse.
Results and Analyses
We present the results of 3D-QA experiments in Sec. 5.2 and the results of OV-Seg experiments in Sec. 5.2. The analyses are as follows:
[…]
    As shown in Sec. 5.2, our model achieves state-of-the-art results on both benchmarks, outperforming recent strong pre-training-based baselines like 3D-VisTA and 3D-LLM. As SceneVerse currently contains only descriptions of objects and scenes, we believe involving more types of language descriptions (e.g., question-answer pairs, dialogues) is a promising direction for further improving model performance on these downstream tasks.
[…]
    As shown in Sec. 5.2, we observe consistent performance improvement of existing 3D backbone models on this task when pre-trained with SceneVerse data. This result validates that the collected data in SceneVerse can effectively boost the performance of existing models on scene understanding tasks. We further provide results of state-of-the-art 3D models that are pre-trained on SceneVerse on the close-vocabulary 3D semantic segmentation task in the supplementary.
5.4 Ablative Studies and Discussion
In this section, we present further discussions on both the data collected in SceneVerse and the GPS model design. We aim to elucidate the effects of data scaling and show more clearly its effectiveness in 3D scene understanding. Regarding the experimental settings and more results discussion, refer to the supplementary. The following points are specifically discussed in this section:
@@ 5.4 Ablative Studies and Discussion
We conduct ablation studies over the amount of data used while pre-training GPS. We consider the model trained with $\frac{1}{8}$, $\frac{1}{4}$, $\frac{1}{2}$ of SceneVerse to show the effectiveness of data-scaling on model performance in the pre-train and zero-shot transfer settings in ScanRefer and SceneVerse-val. As shown in Fig. 4, we observe consistent performance improvement over the increase of data scale for both settings. We provide additional experiments in the supplementary to show that such scaling effect is not only beneficial for 3D-VL grounding but also for other 3D tasks like semantic segmentation [83, 99].
[Uncaptioned image][Uncaptioned image]
Figure 4: Model performance v.s. data scale. Plots show that models consistently improve in both the pre-train and zero-shot transfer settings on ScanRefer and SceneVerse-val with data scaling-up.
Table 7: Ablation on text data source used in model pre-training. All models are tested on ScanRefer with no additional finetuning.
@@
Table 8: Cross domain transfer results of models pre-trained on real and synthetic datasets. “S3D” stands for Structured3D.
@@
Table 9: Ablation on model design on SceneVerse-val. We use “Obj-lvl”, “Scene-lvl” to denote object and scene alignment loss, and “MLM” for the mask language modeling loss.
Obj-lvl. MLM Scene-lvl. Overall Easy Hard
64.8 75.4 48.7
65.2 77.1 47.4
62.4 73.4 45.8
66.9 77.8 50.3
How is the generated data compared with human-annotated data?
We assess the performance of models trained using various scene-text sources, specifically focusing on their performance in the ScanRefer dataset without additional fine-tuning. As shown in Sec. 5.4, models trained with our template-based generated texts and Large Language Model (LLM)-refined texts show significant improvements over models trained solely on ScanRefer. More importantly, these variants of our model already achieve state-of-the-art results compared with previous baselines. This indicates the effectiveness of our text-generation pipeline. Finally, we observe that adding human-annotated data is still beneficial for model performance. However, the improvement is relatively marginal over models trained on our generated data.
What is the role of the synthetic scenes in this scale-up process?
With synthetic data providing large-scale and diverse scene data for 3D-VL tasks, we evaluate the models’ domain transfer (Sim2Real) capability. Specifically, we compare models trained on all real scenes in SceneVerse against models trained exclusively on two synthetic subsets of SceneVerse, i.e., Structured3D and ProcTHOR. As shown in Sec. 5.4, models trained on synthetic subsets demonstrate remarkable performance on their corresponding test sets while suffering when transferred to real or other synthetic scenes. In contrast, the model trained on real scene-text pairs exhibits less performance drop when generalizing to synthetic scenes. This result affirms the domain gap between real and synthetic scenes in 3D-VL grounding and shows that a simple scale-up in the number of scenes is insufficient when naturalness can not be guaranteed. Considering the scalability of our language generation pipeline and the scaling effect shown in our experiments, the rate-determining step for further scaling-up 3D-VL comes to the collection of diverse, high-quality, and realistic scenes that capture natural 3D scene distributions.
How important is the design of each module in GPS?
We provide ablative analyses of our multi-level contrastive alignment design in Sec. 5.4. We mainly consider removing objectives in our model to reveal the effectiveness of each level of alignment. We choose the referral-object-level alignment objective as the default setting and consider removing: (i) object-level alignment objective, (ii) masked language modeling objective, and (iii) scene-level alignment objective. When removing the object-level alignment objective, we learn the object point cloud encoder with the referral-object-level alignment and without pre-training. As shown in Sec. 5.4, we test different models on the SceneVerse-val without additional fine-tuning. Results show that the scene-level alignment objective is crucial for referral object grounding in SceneVerse-val with the $\sim$5performance drop. Similar observations could be made for the model trained without object-level alignment ($\sim$2drop) and masked language modeling objective ($\sim$1.5drop). These results affirm the effectiveness of our model design.
@@ 6 Conclusion
In this work, we scale up 3D-VL for grounded scene understanding. We present SceneVerse, a million-scale 3D-VL dataset covering various scenes and multi-level scene descriptions sourced from both human annotation and our proposed scene-text generation approach. Utilizing SceneVerse, we propose Grounded Pre-training for Scenes (GPS), a model trained with multi-level scene-language contrastive alignment. Through extensive experiments, we show that GPS achieves state-of-the-art results on common 3D-VL tasks including grounding and question answering. We further conduct zero-shot transfer experiments to show the improved generalization performances of GPS trained on SceneVerse compared with previous baselines. We also demonstrate that the scaling effect of SceneVerse is generally beneficial for existing 3D models on 3D-VL tasks like semantic segmentation. We hope our efforts and successful scale-up attempts in SceneVerse could pave the way for new research paradigms in 3D-VL.
Acknowledgement
The authors would like to thank Yaowei Zhang (BIGAI) for his help on online visualization and other colleagues from BIGAI General Vision Lab for fruitful discussions. The authors would also like to thank the anonymous reviewers for their constructive feedback.
@@ Acknowledgement
Appendix 0.A The SceneVerse Dataset
0.A.1 3D Scenes
@@ Appendix 0.A The SceneVerse Dataset
To mitigate the data disparities arising from diverse capture devices across various data sources, we subsample each point cloud to a maximum of $240,000$ points. Each point cloud then undergoes a transformation centered on the central point on the floor, followed by rotation to align the room layout with the axis following the approach by Chen et al. [18].
@@
0.A.2 3D Scene Graph Construction
@@
In-contact vertical supported by embedded into
placed in inside
Non-contact vertical hanging on affixed on
mounted on above
higher than below
@@
Horizontal near(far) to the left of near(far) to the right of
is behind is in front of
close to adjacent to
besides next to
Multi-object between aligned
@@
Scene Graph Construction
Due to the inherent noise and incompleteness in the point cloud representation, automatically extracting precise and comprehensive relationships from the point clouds is a non-trivial task. Below we detail our 3D scene graph construction process, as outlined in Algorithm 1.
@@
In our constructed 3D scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the nodes $\mathcal{V}$ comprises the union of node sets $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level. The hierarchies are determined by the support relationship; for instance, objects supported by the floor constitute $\mathcal{V}_{0}$, while objects supported by the table will form $\mathcal{V}_{1}$, etc. Note that edges originating from one node $v\in\mathcal{V}_{k}$ may only terminate in nearby hierarchies $\mathcal{V}_{k}\cup\mathcal{V}_{k+1}\cup\mathcal{V}_{k+1}$. In other words, edges in the scene graph exclusively connect nodes within the same hierarchical level, or one level higher or lower.
@@
12:     Create node $v_{i}\in\mathcal{V}$ using the centroid $p_{i}$ and bounding box size $b_{i}$ of object
point cloud $P_{i}$
@@
27:        Add in-contact vertical relationship triplets $(v_{i},v_{j},e_{i,j})$ with $\mathrm{RelsType}_{v}$
to $\mathcal{G}$
@@
312:           Add non-contact vertical relationship triplets $(v_{i},v_{k},e_{i,k})$ with
$\mathrm{RelsType}_{v}$ to $\mathcal{G}$
@@
417:     let ${\{v_{i_{1}},v_{i_{2}},...,v_{i_{N}}\}}$ be the $N$ different nodes with the same in-contact
vertical parent node $v_{i}$
@@
0.A.3 Language Generation Details
0.A.3.1 Object Captioning Pipeline
Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features. The detailed object captioning pipeline is outlined in Algorithm 2. Given the multi-view images $\{I_{1},I_{2},\dots,I_{n}\}$, we utilize the point cloud $P_{o}$ of the object $o$ to get the visible points $P_{o,v}^{vis}$ in the images $v$ through rendering. The occlusion score $s_{o,v}^{occ}$ is calculated as the ratio between the number of visible points and the object point cloud. The image is then cropped with the rendered bounding box and processed through BLIP2 [58] to generate the initial object caption $C_{o,v}$. For each initial caption, we calculate its CLIP [78] similarity score between the text and the cropped image, denoted by $s_{o,v}^{clip}$. To get a refined object caption, we select the top $10$ initial captions with the highest CLIP score and minimal occlusion. The selected sentences are fed into a LLM to obtain a coherent summary of the object captions. In this process, we explicitly instruct the language model to identify and correct the potential errors.